OpenStudy (decentnabeel):

what are two numbers that add to 60 and multiply to -500

OpenStudy (decentnabeel):

@mathmale

OpenStudy (decentnabeel):

@rishavraj

OpenStudy (misty1212):

HI!!

OpenStudy (nottim):

Set x+y= 60, and xy=-500. Rearrange and substitute.

OpenStudy (misty1212):

you can probably do this by solving \[x^2+60x-500=0\]

OpenStudy (decentnabeel):

@misty1212 that is not nature number after solving the Euation

OpenStudy (misty1212):

ok then try @NotTim way

OpenStudy (seratul):

You can use the quadratic formula as well.

OpenStudy (decentnabeel):

\[x=10\left(\sqrt{14}-3\right),\:x=-10\left(3+\sqrt{14}\right)\] After Solveing the Rquation that is the Answer come

OpenStudy (misty1212):

\[x+y=60\iff y=60-x\] \[xy=-500\\ x(60-x)=-500\]

OpenStudy (decentnabeel):

@Seratul YEs i have also Solve the Quadratic formula

OpenStudy (misty1212):

pretty sure you get the same equation

OpenStudy (decentnabeel):

yes

OpenStudy (seratul):

Finish solving.

OpenStudy (misty1212):

no one said the answers had to be integers right?

OpenStudy (decentnabeel):

I don't know the answer

satellite73 (satellite73):

did you get to here \[x(60-x)=-500\]?

OpenStudy (seratul):

What did you get after applying the quadratic formula?

OpenStudy (decentnabeel):

yes but answer is not integer x(60-x)=-500 After solveing the answer is.. \[x=-10\left(\sqrt{14}-3\right),\:x=10\left(3+\sqrt{14}\right)\]

OpenStudy (misty1212):

well, they certainly add to 60!!

OpenStudy (misty1212):

not sure why you are certain the answers are whole numbers in general, if you solve a quadratic equation, the answers have radicals

OpenStudy (decentnabeel):

@Seratul \[x=-10\left(\sqrt{14}-3\right),\:x=10\left(3+\sqrt{14}\right)\] that is answer after applying Quadratic formula

satellite73 (satellite73):

or \[30-10\sqrt{14}, 30+10\sqrt{14}\]

OpenStudy (seratul):

Then that's your answer. You can continue to solve it if you like to what satellite said.

satellite73 (satellite73):

as @misty1212 said, they add up to 60 for certain you can check to see if they multiply to get \(-500\)